Find The Sum or Difference.
22. -3 1/4 +(-2 2/5)
Any help with this equation would be GREATLY appreciated.

Answer by edjones(8007)

-3 1/4 +(-2 2/5)
=-13/4 - 12/5
=-65/20- 48/20 LCD=20
= -113/20
= -5 13/20
Ed